# SC-GS: Sparse-Controlled Gaussian Splatting for Editable Dynamic Scenes

### v2

@@ Abstract
Novel view synthesis for dynamic scenes is still a challenging problem in computer vision and graphics. Recently, Gaussian splatting has emerged as a robust technique to represent static scenes and enable high-quality and real-time novel view synthesis. Building upon this technique, we propose a new representation that explicitly decomposes the motion and appearance of dynamic scenes into sparse control points and dense Gaussians, respectively. Our key idea is to use sparse control points, significantly fewer in number than the Gaussians, to learn compact 6 DoF transformation bases, which can be locally interpolated through learned interpolation weights to yield the motion field of 3D Gaussians. We employ a deformation MLP to predict time-varying 6 DoF transformations for each control point, which reduces learning complexities, enhances learning abilities, and facilitates obtaining temporal and spatial coherent motion patterns. Then, we jointly learn the 3D Gaussians, the canonical space locations of control points, and the deformation MLP to reconstruct the appearance, geometry, and dynamics of 3D scenes. During learning, the location and number of control points are adaptively adjusted to accommodate varying motion complexities in different regions, and an ARAP loss following the principle of as rigid as possible is developed to enforce spatial continuity and local rigidity of learned motions. Finally, thanks to the explicit sparse motion representation and its decomposition from appearance, our method can enable user-controlled motion editing while retaining high-fidelity appearances. Extensive experiments demonstrate that our approach outperforms existing approaches on novel view synthesis with a high rendering speed and enables novel appearance-preserved motion editing applications. Project page: https://yihua7.github.io/SC-GS-web/
@@ 1 Introduction
Novel view synthesis from a monocular video is a crucial problem with many applications in virtual reality, gaming, and the movie industry. However, extracting scene geometry and appearance from limited observations  [23, 41, 24] is challenging. Concurrently, real-world scenes often contain dynamic objects, which pose additional challenges in representing object movements accurately to reflect real-world dynamics [30, 15, 26, 27, 14].
@@
Recent advancements in this area are primarily driven by neural radiance fields (NeRF)  [23, 1, 41, 2, 3], which utilizes an implicit function to simultaneously learn scene geometry and appearance from multi-view images. Various approaches have been proposed to model dynamic objects by predicting scene flows [30, 15, 43, 38, 6, 26, 27, 11, 16] leveraging plane- or volume grid-based representations [31, 4, 33, 20, 7] and temporal interpolation [14, 28]. Despite significant progress, NeRF-based representations still struggle with low rendering qualities, speeds, and high memory usage. This issue is particularly evident when rendering at high resolutions [50, 8, 24], as they necessitate sampling hundreds of query points along each ray to predict color and opacity for rendering a single pixel.
@@
Most recently, Gaussian splatting [12], which represents a scene as 3D Gaussians, has shown remarkable performance in terms of rendering quality, resolution, and speed. Utilizing a point-based scene representation, this method rasterizes 3D Gaussians to render images from specified views. It enables fast model training and real-time inference, achieving state-of-the-art (SOTA) visual quality. However, its existing formulation only applies to static scenes. It remains a challenge to incorporate object motion into the Gaussian representation without compromising rendering quality and speed. An intuitive method [47] involves learning a flow vector for each 3D Gaussian. While this method can yield high-quality rendering results, it incurs a significant time cost for training and inference. Moreover, modeling motion in such a high-dimensional space leads to noisy trajectories and poor generalization in novel views, as demonstrated in Fig. 6 (a).
@@
Motivated by the observation that real-world motions are often sparse, spatially continuous, and locally rigid, we propose an approach to enrich Gaussian splatting with learnable sparse control points ($\approx$512) compared to the number of Gaussians ($\approx$100K), in a much more compact space for modeling scene dynamics. These control points are associated with time-varying 6 DoF transformations parameterized as rotation using quaternion and translation parameters, which can be locally interpolated through learned interpolation weights to yield the motion field of dense Gaussians. Instead of directly optimizing 6 DoF parameters for each control point, we utilize an MLP conditioned on time and location to predict them. This strategy reduces learning complexities, enhances model capabilities, and provides motions with improved spatial and temporal continuity and coherence. Then, we jointly learn the canonical space 3D Gaussian parameters, locations, and radius of sparse control points at canonical space and the MLP to predict their time-varying transformations from a monocular video for dynamic novel view synthesis. During learning, we introduce an adaptive strategy to adaptively change the number of sparse points to accommodate motion complexities in different regions and employ an ARAP loss that encourages the learned motions to be locally as rigid as possible.
@@ 2 Related Work
Dynamic NeRF. Novel view synthesis has been a prominent topic in the academic field for several years. NeRF [23] models static scenes implicitly with MLPs, and many works [30, 15, 43, 38, 26, 27, 11] have expanded its usage to dynamic scenes via a deformation field. Some methods [9, 14, 28] represent dynamic scenes as 4D radiance fields but face extensive computational costs due to ray point sampling and volume rendering. Several acceleration approaches have been used for dynamic scene modeling. DeVRF [20] introduces a grid representation, and IBR-based methods [18, 16, 17, 45] use multi-camera information for quality and efficiency. Other methods used primitives [21], predicted MLP maps [29], or grid/plane-based structures [33, 7, 4, 31, 39, 40] for speed and performance in various dynamic scenes. However, hybrid models under perform with high-rank dynamic scenes due to their low-rank assumption.
Dynamic Gaussian Splatting. Gaussian Splatting [12] offers improved rendering quality and speed for radiance fields. Several concurrent works have adapted 3D Gaussians for dynamic scenes. Luiten et al. [22] utilizes frame-by-frame training, suitable for multi-view scenes. Yang et al. [47] separate scenes into 3D Gaussians and a deformation field for monocular scenes but face slow training due to an extra MLP for learning Gaussian offsets. Following [47], Wu et al. [42] replaced the MLP with multi-resolution hex-planes [4] and a lightweight MLP. Yang et al. [48] include time as an additional feature in 4D Gaussians but face quality issues compared to constraints in canonical space. Our work proposes using sparse control points to drive the deformation of 3D Gaussians, which enhances rendering quality and reduces MLP query overhead. The learned control point graph can also be used for motion editing.
3D Deformation and Editing. Traditional deformation methods in computer graphics are typically based on Laplacian coordinates [19, 36, 35, 34], Poisson equation [51] and cage-based approaches [49, 54]. These methods primarily focus on preserving the geometric details of 3D objects during the deformation process. In recent years, there have been other approaches [52, 44] that aim to edit the scene geometry learned from 2D images. These methods prioritize the rendering quality of the edited scene. Our approach falls into this category. However, instead of relying on the implicit and computationally expensive NeRF-based approach, our method employs an explicit point-based control graph deformation strategy and Gaussian rendering, which is more intuitive and efficient.
@@ 6.2 Quantitative Comparisons
Reconstructed Motion
Edited Motion
@@ 6.5 Motion Editing
7 Conclusion and Limitations
We present a method driving 3D Gaussians using control points and a deformation MLP, learnable from dynamic scenes. Our approach, combining a compact motion representation with adaptive learning strategies and rigid constraints, allows high-quality dynamic scene reconstruction and motion editing. Experiments showed our method outperforms existing approaches in the visual quality of synthesized dynamic novel views. However, limitations exist. The performance is prone to inaccurate camera poses, leading to reconstruction failures on datasets with inaccurate poses such as HyperNeRF [27]. The approach can also overfit datasets with sparse viewpoints or limited angle coverage. Finally, the method has been tested only in scenes with modest motion changes. Extending its applicability to intense movements remains an area of exploration.

### v3

@@ Abstract
Novel view synthesis for dynamic scenes is still a challenging problem in computer vision and graphics. Recently, Gaussian splatting has emerged as a robust technique to represent static scenes and enable high-quality and real-time novel view synthesis. Building upon this technique, we propose a new representation that explicitly decomposes the motion and appearance of dynamic scenes into sparse control points and dense Gaussians, respectively. Our key idea is to use sparse control points, significantly fewer in number than the Gaussians, to learn compact 6 DoF transformation bases, which can be locally interpolated through learned interpolation weights to yield the motion field of 3D Gaussians. We employ a deformation MLP to predict time-varying 6 DoF transformations for each control point, which reduces learning complexities, enhances learning abilities, and facilitates obtaining temporal and spatial coherent motion patterns. Then, we jointly learn the 3D Gaussians, the canonical space locations of control points, and the deformation MLP to reconstruct the appearance, geometry, and dynamics of 3D scenes. During learning, the location and number of control points are adaptively adjusted to accommodate varying motion complexities in different regions, and an ARAP loss following the principle of as rigid as possible is developed to enforce spatial continuity and local rigidity of learned motions. Finally, thanks to the explicit sparse motion representation and its decomposition from appearance, our method can enable user-controlled motion editing while retaining high-fidelity appearances. Extensive experiments demonstrate that our approach outperforms existing approaches on novel view synthesis with a high rendering speed and enables novel appearance-preserved motion editing applications.
@@ 1 Introduction
Novel view synthesis from a monocular video is a crucial problem with many applications in virtual reality, gaming, and the movie industry. However, extracting scene geometry and appearance from limited observations  [30, 49, 31] is challenging. Concurrently, real-world scenes often contain dynamic objects, which pose additional challenges in representing object movements accurately to reflect real-world dynamics [37, 19, 33, 34, 18]. Recent advancements in this area are primarily driven by neural radiance fields (NeRF)  [30, 37, 19, 66], which utilizes an implicit function to simultaneously learn scene geometry [29, 26] and textures [12, 57] from multi-view images. Despite significant progress, NeRF-based representations still struggle with low rendering speeds and high memory usage. This issue is particularly evident when rendering at high resolutions [62, 6, 31], as they necessitate sampling hundreds of query points along each ray to predict color and opacity.
@@
Most recently, Gaussian splatting [13] has shown remarkable performance in terms of rendering quality, resolution, and speed. Utilizing a point-based [14, 53, 2, 67, 15, 10, 46] scene representation, this method rasterizes 3D Gaussians to render images from specified views. It enables fast model training and real-time inference, achieving state-of-the-art (SOTA) visual quality. However, its existing formulation only applies to static scenes. It remains a challenge to incorporate object motion into the Gaussian representation without compromising rendering quality and speed. An intuitive method is to learn a flow vector for each 3D Gaussian. However, it will incur a significant time cost for training and inference. Moreover, it also leads to noisy trajectories and poor generalization in novel views, as demonstrated in Fig. 6 (a).
@@
Motivated by the observation that real-world motions are often sparse, spatially continuous, and locally rigid, we propose to drive 3D Gaussians with learnable sparse control points ($\approx$512) compared to the number of Gaussians ($\approx$100K), in a much more compact space for modeling scene dynamics. These control points are associated with time-varying 6 DoF transformations parameterized as rotation using quaternion and translation parameters, which can be locally interpolated through learned interpolation weights to yield the motion field of dense Gaussians. These 6 DoF parameters on control points are predicted by an introduced MLP conditioned on time and location. Then, we jointly learn the canonical space 3D Gaussian parameters, locations, and radius of sparse control points at canonical space and the MLP for dynamic novel view synthesis. During learning, we introduce an adaptive strategy to adaptively change the number of sparse points to accommodate motion complexities in different regions and employ an ARAP loss that encourages the learned motions to be locally rigid.
@@ 2 Related Work
Dynamic NeRF. Novel view synthesis has been a prominent topic in the academic field for several years. NeRF [30] models static scenes implicitly with MLPs, and many works [37, 19, 52, 45, 33, 34, 11, 66] have expanded its usage to dynamic scenes via a deformation field. Some methods [7, 18, 35] represent dynamic scenes as 4D radiance fields but face extensive computational costs due to ray point sampling and volume rendering. Several acceleration approaches have been used for dynamic scene modeling. DeVRF [25] introduces a grid representation, and IBR-based methods [23, 20, 22, 55] use multi-camera information for quality and efficiency. Other methods used primitives [27], predicted MLP maps [36], or grid/plane-based structures [40, 5, 1, 38, 47, 48] for speed and performance in various dynamic scenes. However, hybrid models underperform with high-rank dynamic scenes due to their low-rank assumption.
Dynamic Gaussian Splatting. Gaussian Splatting [13, 51] offers improved rendering quality and speed for radiance fields. Several concurrent works have adapted 3D Gaussians for dynamic scenes. Luiten et al. [28] utilizes frame-by-frame training, suitable for multi-view scenes. Yang et al. [58] separate scenes into 3D Gaussians and a deformation field for monocular scenes but face slow training due to an extra MLP for learning Gaussian offsets. Following [58], Wu et al. [50] replaced the MLP with multi-resolution hex-planes [1] and a lightweight MLP. Yang et al. [59] include time as an additional feature in 4D Gaussians but face quality issues compared to constraints in canonical space. Our work proposes using sparse control points to drive the deformation of 3D Gaussians, which enhances rendering quality and reduces MLP query overhead. The learned control point graph can also be used for motion editing.
3D Deformation and Editing. Traditional deformation methods in computer graphics are typically based on Laplacian coordinates [24, 43, 42, 41, 8], Poisson equation [63] and cage-based approaches [61, 69]. These methods primarily focus on preserving the geometric details of 3D objects during the deformation process. In recent years, there have been other approaches [64, 65, 54, 70, 21] that aim to edit the scene geometry learned from 2D images. These methods prioritize the rendering quality of the edited scene. Our approach falls into this category. However, instead of relying on the implicit and computationally expensive NeRF-based approach, our method employs an explicit point-based control graph deformation strategy and Gaussian rendering, which is more intuitive and efficient.
@@ 6.2 Quantitative Comparisons
Reconstructed Motion Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Edited Motion Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 6.5 Motion Editing
7 Conclusion and Future Works
We present a method driving 3D Gaussians using control points and a deformation MLP, learnable from dynamic scenes. Our approach, combining a compact motion representation with adaptive learning strategies and rigid constraints, allows high-quality dynamic scene reconstruction and motion editing. Experiments showed our method outperforms existing approaches in the visual quality of synthesized dynamic novel views. However, limitations exist. The performance is prone to inaccurate camera poses, leading to reconstruction failures on datasets with inaccurate poses such as HyperNeRF [34]. The current approach also faces limitations in handling common specular effects, resulting in limited improvement on NeRF-DS [56] datasets. To address this, future work could focus on extending the method by incorporating Spec-Gaussian [60] with a specialized specular design. This enhancement would enable more accurate modeling of highlight and mirror effects. Furthermore, the presence of blurriness in videos with dynamic objects should be considered. To enhance the robustness of the proposed method, incorporating deblurring techniques [3, 16] for novel view synthesis can address this issue effectively.
Acknowledgement
This work has been supported by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422), and RGC Matching Fund Scheme (RMGS). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.